# Formation of Bound States in Quintessence Alternative Theories

### v1

@@ Abstract
We study the formation and behaviour of bound states formed outside the horizon of a black hole in the presence of quintessence matter. Calculating the Regge and Wheeler potential for general metric function, we find that the presence of quintessence influences significantly the metric function and the Hawking temperature. We show that large black holes radiate less in the presence of large quintessence matter and it seems to live longer, while small black holes radiate more.
@@ I Introduction
Bound states for particles around black holes is a very interesting effect and a detailed study was carried out in Grain:2007gn . A similar study can be done by the investigation of QNMs. In both cases the goal is to find the characteristic complex frequencies for a scalar field which propagates in a black hole background. The constraints which the scalar field has to satisfy are boundary conditions at the black hole horizon and at spatial infinity. However, the physical meanings of bound states and quasi-normal modes are very different and correspond to different energy ranges. In Grain:2007gn a systematic study was carried out of quanta trapped between the finite black hole potential barrier and the infinitely thick well which works as a barrier preventing the particles from reaching infinity. These bound states may have possible cosmological consequences related to the primordial power spectrum barrau3 ; Carr:2005zd . Also the phenomenology of these bound states and their spectra should be studied in details and experiments associated with light black holes should also take into account those states.
@@
One basic tool to examine the stability of a black hole is to study the Regge and Wheeler potential. Regge and Wheeler regwh first described a decomposition of perturbations on a spherically symmetric background geometry into tensor harmonics and wrote the perturbation equations in terms of these harmonics. They found that it is possible to put the equations for the oscillatory motion into the form of a single second-order Schrodinger-type differential equation with an effective potential. Then the study of these equations can give us important information on the stability of the black hole The first metric that was studied by Regge and Wheeler was the Schwarzschild metric.
@@
The formation and propagation of quantum bound states in the vicinity of a Galileon black hole was discussed in Koutsoumbas:2018gbd . The Galileon black holes are generated by a scalar field coupled to the Curvature Kolyvaris:2011fk ; Rinaldi:2012vy ; Kolyvaris:2013zfa and they are spherically symmetric hairy black holes. Classically this coupling has been studied extensively; it strongly influences the propagation of the scalar field compared to a scalar field minimally coupled to gravity. Also the stability of the background Galileon black hole was studied and it was found that bound states were trapped in a Regge and Wheeler potential well or penetrating the horizon of the Galileon black hole depending on the strength of the derivative coupling of a scalar field to curvature. The formation of quantum bound states requires that the Klein-Gordon equation in a black hole background has to exhibit a radial potential containing a local well for given ranges of black hole parameters.
@@
Recent astronomical observations suggest that our Universe is expanding at an accelerating rate Riess1 ; Perlmutter1999 . Responsible for this expansion is the dark sector of our Universe consisting mainly of dark energy, while the presence of a cosmological constant cannot explain this expansion Carroll2001 ; Copeland2006 . Approximately 70% of the energy density in the universe is in the form of an exotic, negative-pressure component, dubbed dark energy. The observational bounds on the properties of the dark energy have continued to tighten. Taking $w$ to be the ratio of pressure to density for the dark energy $w=p_{DE}/\rho_{DE},$ recent observational constraints the range of $-1\leq w\leq-\frac{1}{3}$, which causes the acceleration. The border case of $w=-1$ of the extraordinary quintessence covers the cosmological constant term. As it was recognized Suss ; B ; W , the acceleration represents a challenge for the consistent theory of quantum gravity, because of the outer horizon, which does not allow one to introduce the observable S-matrix in terms of asymptotic past and future states. The outer horizon of de Sitter space significantly differs from the inner horizon of black hole, which has asymptotically flat space far away from the black hole.
@@
The presence of a quintessence field to explain the Universe expansion modifies General Relativity and new modified Gravity theories were proposed with the presence of dynamical scalar fields Hellerman2001 . In these theories black hole solutions were found in the presence of the quintessence matter field. One of the first black hole solution in the presence of quintessence was discussed in Kiselev:2002dx . New static spherically-symmetric exact solutions of Einstein equations with the quintessential matter surrounding a black hole charged or uncharged were presented. A condition of additivity and linearity in the energy-momentum tensor was found which allows to get correct limits to the known solutions for the electromagnetic static field. Further studies in the presence of quintessence matter were carried out and new black holes solutions were obtained, and the impacts of the quintessence field on their thermal quantities, shadows, and quasinormal modes were discussed Chen2008 -BB5 . The presence of dark matter and quintessence fields outside the horizon of a black hole has an impact on their thermodynamic properties, such as the temperature, and introduces new critical phenomena. These modifications could lead to generalizations of classical thermodynamic results making this area of study very interesting leading to connection between a black hole event horizon and its entropy Bekenstein1973 ; Hawking1974 .
@@
The aim of this work is to study the formation and behaviour of bound states formed outside the horizon of a hairy black hole in which the hair is formed by a quintessence matter. Then considering a general form of the metric function we derive a Schrödinger equation in which the Regge and Wheeler potential appears. Considering also the thermodynamics of the solution we discuss the formation and the behaviour of the bound states formed for various values of the quintessence parameter w.
@@
The work is organized as follows. In section II we review the Regge-Wheeler potential for a fairly general metric function. In section III the quintessence is introduced and some aspects of its thermodynamics are described. Section IV deals with the Regge-Wheeler potential for scalar excitations, while in section V mass eigenstates are calculated. Section VI contains our conclusions.
@@ II Generalized Regge-Wheeler Potential
We decompose the various fields into radial, spherical and time parts, using the ansatz
@@ III Quintessence Matter and its Thermodynamics
In this section we will introduce quintessence matter outside the horizon of a black hole and we study the thermodynamics of the coupling system and discuss the formation of bound states.
@@ III.1 Description of the Quintessence
As we have already discussed quintessence is a scalar field, whose equation of state parameter $w$ is defined as the ratio of its pressure $p$ and its energy density $\rho$ which is given by a kinetic term and a potential term as
$w=\frac{p}{\rho}=\frac{\frac{1}{2}\dot{\phi}^{2}-V(\phi)}{\frac{1}{2}\dot{\phi}^{2}+V(\phi)},\ \ -1\leq w\leq-\frac{1}{3}~{}.$ (19)
The energy-momentum tensor in the presence of quintessence is given by Kiselev:2002dx :
$T_{t}^{t}=\rho,\ T_{k}^{l}=3w\rho\left[-(1+3B)\frac{r_{k}r^{l}}{r^{2}}+B\delta_{k}^{l}\right],\ B=-\frac{1+3w}{6w}~{}.$ (20)
Averaging over angles, taking into account that $<r_{k}r^{l}>=\frac{1}{3}r^{2}\delta_{k}^{l},$ yields
$<T_{k}^{l}>=-p\delta_{k}^{l}~{}.$ (21)
@@
the Einstein equations read:
@@
There exist two solutions, the well-known Schwarzschild solution $h=-\frac{2M}{r}$ and a novel one
$h=\frac{a}{r^{1+3w}}~{}.$ (24)
The Einstein equations yield:
@@
Notice that $R$ is singular at $r\to 0,$ except when $w=0,\ \frac{1}{3},\ -1.$ Additivity and linearity yield the final solution for the metric: $f=1-\frac{2M}{r}-\frac{a}{r^{1+3w}}.$ The generalized solution to the Reissner–Nordström–anti-deSitter black hole reads:
@@
As we can see the metric function in the AdS space apart from the usual parameters of mass and charge because of the presence of quintessence matter also depends on two more parameters $a$ and $w,$ which, as we will discuss, play a crucial role in the thermodynamics of the system and the formation of bound states.
@@ III.2 Thermodynamics in AdS Spacetime
In Fig. 1, upper left panel, we plot the metric function for $Q=0.00,\ w=-0.80$ and various values of the $a$ parameter. In the upper right panel of the same figure, we plot the metric function for $Q=0.00,\ a=0.80$ and various values of the $w$ parameter. In the panel down left, we plot the metric function for $Q=0.20,\ w=-0.80$ and various values of the $a$ parameter. Finally, in the panel down right, we plot the metric function for $Q=0.20,\ a=0.80$ and various values of the $w$ parameter. As we can see, the existence of a non-zero charge $Q$ in the black hole tends to decrease the value of the metric function. The qualitative characteristics of the graphs do not change much with respect to the $Q=0.00$ cases.
@@
Figure 1: (a) Metric function $g\equiv f(r)$ versus $r$ for $Q=0.00,$ $w=-0.80$ and $a=0.20,$ (blue curve) $a=0.40$ (purple curve) $a=0.60$ (brown curve) and $a=0.80$ (cyan curve). (b) Metric function $g\equiv f(r)$ versus $r$ for $Q=0.00,$ $a=0.80$ and $w=-0.40,$ (blue curve) $w=-0.60$ (purple curve) $w=-0.80$ (brown curve) and $w=-1.00$ (cyan curve). (c) Metric function $g\equiv f(r)$ versus $r$ for $Q=0.20,$ $w=-0.80$ and $a=0.20,$ (blue curve) $a=0.40$ (purple curve) $a=0.60$ (brown curve) and $a=0.80$ (cyan curve). (d) Metric function $g\equiv f(r)$ versus $r$ for $Q=0.20,$ $a=0.80$ and $w=-0.40,$ (blue curve) $w=-0.60$ (purple curve) $w=-0.80$ (brown curve) and $w=-1.00$ (cyan curve).
In Fig. 2, upper left panel, we plot the temperature versus $r_{+}$ for $Q=0.00,\ w=-0.80$ and various values of the $a$ parameter. In the upper right panel of the same figure, we plot the temperature versus $r_{+}$ for $Q=0.00,\ a=0.80$ and various values of the $w$ parameter. In the panel down left, we plot the temperature versus $r_{+}$ for $Q=0.20,\ w=-0.80$ and various values of the $a$ parameter. Finally, in the panel down right, we plot the temperature versus $r_{+}$ for $Q=0.20,\ a=0.80$ and various values of the $w$ parameter. As we can see, the existence of a non-zero charge $Q$ in the black hole tends to decrease the value of the metric function.
The dependence of the temperature on $a$ is monotonous, making it smaller for increasing $a.$ This characteristic does not survive at $Q=0.20,$ as is obvious from the left panels of the figure.
@@
The $w$ dependence is less trivial: it is smaller for large $r_{+},$ while for modest $r_{+}$ it is larger. For $Q=0.00,$ at some horizon size, the temperature has a minimum. The temperature of the black hole becomes smaller as $|w|$ increases from $\frac{1}{3}$ to $1$ for large enough black holes (big $r_{+}$), while it becomes larger for small $r_{+}.$ Thus large black holes radiate less for large $|w|,$ while small black holes radiate more. Thus large black holes seem to be longer lived for large $|w|.$
@@
Figure 2: (a) Temperature versus $r_{h}\equiv r_{+}$ for $Q=0.00,$ $w=-0.80$ and $a=0.20,$ (blue curve) $a=0.40$ (purple curve) $a=0.60$ (brown curve) and $a=0.80$ (khaki curve). (b) Temperature versus $r_{h}\equiv r_{+}$ for $Q=0.00,$ $a=0.80$ and $w=-0.40,$ (blue curve) $w=-0.60$ (purple curve) $w=-0.80$ (cyan curve) and $w=-1.00$ (cyan curve). (c) Temperature versus $r_{h}\equiv r_{+}$ for $Q=0.00,$ $w=-0.80$ and $a=0.20,$ (blue curve) $a=0.40$ (purple curve) $a=0.60$ (brown curve) and $a=0.80$ (cyan curve). (d) Temperature versus $r_{h}\equiv r_{+}$ for $Q=0.00,$ $a=0.80$ and $w=-0.40,$ (blue curve) $w=-0.60$ (purple curve) $w=-0.80$ (brown curve) and $w=-1.00$ (cyan curve).
@@ IV Regge-Wheeler potentials
We concentrate specifically on the potential for scalar and gravitational perturbations. The potential with zero spin $V_{0}$ for scalar perturbations read
@@
We start in Fig. 3, left panel, with the Regge-Wheeler potentials, when $a=0,$ so that quintessence plays no role at all; we set $M=0.05$ and let $Q$ take the values 0.00, 0.04, 0.070.000.040.070.00,\ 0.04,\ 0.070.00 , 0.04 , 0.07 and $0.10$ It is obvious that, as $Q$ increases, the left barrier of the potential becomes smaller, and at $Q=0.10$ it disappears completely. In this case it is obvious that the potential cannot support bound states: any scalar excitation will end into the black hole horizon. This characteristic is independent from the role of quintessence. We continue in the right panel with the case $a=0.8,\ w=-0.80$ and $M=0.05$ and the same values for the charge $Q$ as before. The picture is qualitatively the same, but the presence of the quintessence tends to reduce the size of the potentials.
@@
Figure 3: Left panel: Potentials for $a=0,\ l=0,\ M=0.05$ and four values of the charge: $Q=0.00$ (pink curve), $0.04$ (blue curve), $0.07$ (khaki curve) and $0.10$ (green curve). Right panel: Potentials for $a=0.8,\ w=-0.80,\ M=0.05$ and $Q=0.00$ (pink curve), $0.04$ (blue curve), $0.07$ (khaki curve) and $0.10$ (green curve).
@@
We go ahead in Fig. 4 with the potentials for four different values of $w,$ namely $-0.40,\ -0.60,\ -0.80,\ -1.00,$ while $l=0,\ M=0.05,$ for $Q=0.00$ (left panel) and $Q=0.04$ (right panel). We observe that, for both values of the charge, the potential for $w=-0.40$ does not support bound states, but a left barrier gradually develops, as $|w|$ increases and bound states become possible.
@@
Figure 4: Left panel: Potentials for $a=0.8,\ l=0,\ M=0.05,\ Q=0.00,$ and four values for the parameter $w:$ $w=-0.40$ (blue curve), $-0.60$ (pink curve), $-0.80$ (khaki curve) and $-1.00$ (green curve). Right panel: Same as in the left panel, but $Q=0.04.$
[…]
@@
Figure 5: Left panel: Potentials for $a=0.8,\ M=0.05,\ Q=0.00,\ w=-0.60$ and three values for the angular momentum: $l=0$ (blue curve), $l=1$ (pink curve), and $l=2$ (green curve). Right panel: Potentials for $l=0,\ M=0.05,\ Q=0.00,\ w=-0.60$ and four values for the parameter $a:$ $a=0.2$ (blue curve), $a=0.4$ (pink curve), $a=0.6$ (khaki curve), and $a=0.8$ (green curve).
V Formation of Bound States

### v2

@@ Abstract
We study the formation and behaviour of bound states formed outside the horizon of a black hole in the presence of quintessence matter. Calculating the Regge and Wheeler potential for general metric function, we find that the presence of quintessence influences significantly the metric function and the Hawking temperature. We show that large black holes radiate less in the presence of quintessence matter and it seems to live longer, while small black holes radiate more in comparison with the model in the absence of quintessence. Bound states emerge at large enough quintessence parameter $|w|$ or angular momentum.
@@ I Introduction
Bound states for particles around black holes is a very interesting effect and a detailed study was carried out in Grain:2007gn . A similar study can be done by the investigation of QNMs. In both cases the goal is to find the characteristic complex frequencies for a scalar field which propagates in a black hole background. The constraints which the scalar field has to satisfy are boundary conditions at the black hole horizon and at spatial infinity. However, the physical meanings of bound states and quasi-normal modes are very different and correspond to different energy ranges. In Grain:2007gn a systematic study was carried out of quanta trapped between the finite adS black hole potential barrier and the infinitely thick well, which works as a barrier preventing the particles from reaching infinity. These bound states may have cosmological consequences related to the primordial power spectrum barrau3 ; Carr:2005zd . Also the phenomenology of these bound states and their spectra should be studied in detail and experiments associated with light black holes should also take into account those states.
@@
One basic tool for the examination of the stability of a black hole is to study the Regge and Wheeler potential. Regge and Wheeler regwh first described a decomposition of perturbations on a spherically symmetric background geometry into tensor harmonics and wrote the perturbation equations in terms of these harmonics. They found that it is possible to put the equations for the oscillatory motion into the form of a single second-order Klein-Gordon-type differential equation with an effective potential. Then the study of these equations can give us important information on the stability of the black hole. The first metric that was studied by Regge and Wheeler was the Schwarzschild metric.
@@
The formation and propagation of quantum bound states in the vicinity of a Galileon black hole was discussed in Koutsoumbas:2018gbd . The Galileon black holes are generated by a scalar field coupled to the Curvature Kolyvaris:2011fk ; Rinaldi:2012vy ; Kolyvaris:2013zfa and they are spherically symmetric hairy black holes. Classically this coupling has been studied extensively; it strongly influences the propagation of the scalar field compared to a scalar field minimally coupled to gravity. Also the stability of the background Galileon black hole was studied and it was found that bound states were trapped in a Regge and Wheeler potential well or penetrating the horizon of the Galileon black hole depending on the strength of the derivative coupling of a scalar field to curvature. The formation of quantum bound states requires that there is a radial potential containing a local well for given ranges of black hole parameters in the KG equation.
@@
Recent astronomical observations suggest that our Universe is expanding at an accelerating rate Riess1 ; Perlmutter1999 . Responsible for this expansion is the dark sector of our Universe consisting mainly of dark energy, while the presence of a cosmological constant cannot explain this expansion Carroll2001 ; Copeland2006 . Approximately 70% of the energy density in the universe is in the form of an exotic, negative pressure component, dubbed dark energy. The observational bounds on the properties of the dark energy have continued to tighten. Taking $w$ to be the ratio of pressure to density for the dark energy $w=p_{DE}/\rho_{DE},$ recent observational constraints the range of $1\leq w\leq-\frac{1}{3}$, which causes the acceleration. The border case of w=1w=-1 of the extraordinary quintessence covers the cosmological constant term. As it was recognized Suss ; B ; W , the acceleration represents a challenge for the consistent theory of quantum gravity, because of the outer horizon, which does not allow one to introduce the observable S-matrix in terms of asymptotic past and future states. The outer horizon of de Sitter space significantly differs from the inner horizon of black hole.
@@
The presence of a quintessence field to explain the Universe expansion modifies General Relativity and new modified Gravity theories were proposed with the presence of dynamical scalar fields Hellerman2001 . In these theories black hole solutions were found in the presence of the quintessence matter field. One of the first black hole solutions in the presence of quintessence was discussed in Kiselev:2002dx . New static spherically-symmetric exact solutions of Einstein equations with the quintessential matter surrounding a black hole charged or uncharged were presented. A condition of additivity and linearity in the energy-momentum tensor was found which allows to get correct limits to the known solutions for the electromagnetic static field. Further studies in the presence of quintessence matter were carried out, new black hole solutions were obtained, and the impacts of the quintessence field on their thermal quantities, shadows, and quasinormal modes were discussed Chen2008 -BB5 . The presence of dark matter and quintessence fields outside the horizon of a black hole has an impact on their thermodynamic properties, such as the temperature, and introduces new critical phenomena. These modifications could lead to generalizations of classical thermodynamic results making this area of study very interesting leading to a connection between a black hole event horizon and its entropy Bekenstein1973 ; Hawking1974 .
@@
Motivated by the work in Kiselev:2002dx in which quintessential matter surrounding a background spherically-symmetric black hole backreacts with it dressing it with quintessential hair, we study the formation and behaviour of bound states formed outside the horizon of this hairy black hole. Then considering a general form of the metric function we derive a Schrödinger equation in which the Regge and Wheeler potential appears. Considering the thermodynamics of the solution we discuss the formation and the behaviour of the bound states formed for various values of the quintessence parameter $w$ for the considered hairy black hole. We also consider matter which does not back-react with the quintessence black hole and we study whether the Regge and Wheeler potential is affected by this matter outside the quintessence black hole.
In summary, we study the formation and the behaviour of bound states outside the horizon of a background spherically-symmetric black hole, which is dressed with quintessential hair. Calculating the Regge and Wheeler potential we study the bound states which are trapped outside the horizon of the black hole depending on the quintessence parameter $w.$ We find that the Regge and Wheeler potential for small $|w|$ does not support bound states. However, as the quintessence parameter $|w|$ increases, formation of bound states becomes possible. In addition, when the angular momentum parameter $l$ increases, more bound states are possible and their imaginary parts are much smaller. We also find that, as the charge $Q$ of the black holes increases, the black hole temperature approaches zero on the horizon, in contrast with the case with $Q=0;$ in addition, less bound states are formed. We work with the black hole presented in Kiselev:2002dx in which quintessential matter surrounding a background spherically-symmetric black hole back-reacts dressing it with quintessential hair. We also discuss the thermodynamic properties of the bound states formed for various values of the quintessence parameter $w.$
@@
The work is organized as follows. In section II we review the Regge-Wheeler potential for a fairly general metric function. In section III the quintessence is introduced and some aspects of its thermodynamics are described. Section IV deals with the Regge-Wheeler potential for scalar excitation, while in section V mass eigenstates are calculated. Section VI contains our conclusions.
@@ II Generalized Regge-Wheeler Potential
We decompose the various fields into radial, angular and time parts, using the ansatz
@@ III Quintessence Matter and its Thermodynamics
In this section we will introduce quintessence matter outside the horizon of a black hole, we study the thermodynamics of the coupled system and discuss the formation of bound states.
@@ III.1 Description of the Quintessence
As we have already discussed, quintessence is a scalar field, whose equation of state parameter $w$ is defined as the ratio of its pressure $p$ and its energy density $\rho$ and is given by
$w=\frac{p}{\rho},\ \ -1\leq w\leq-\frac{1}{3}~.$ (19)
@@
the components of the energy-momentum tensor Kiselev:2002dx read:
@@
There exist two solutions for the Einstein equations, the well-known Schwarzschild solution $h=-\frac{2M}{r}$ and a novel one
$h=\frac{a}{r^{1+3w}},$ (22)
where $a$ and $w$ are constants. The Einstein equations yield:
@@
Notice that $R$ is singular at $r\to 0,$ except when w=0,13,1.w=0,\ $\frac${1}{3},\ -1.
Additivity and linearity yield the final solution for the metric: $f=1-\frac{2M}{r}-\frac{a}{r^{1+3w}}.$ The generalized solution to the Reissner–Nordström–anti-deSitter black hole reads:
@@
The metric function in the AdS Black Hole, apart from the usual dependence mass and charge, it also depends on two more parameters $a$ and $w,$ which play a crucial role in the thermodynamics of the system and the formation of bound states.
@@ III.2 Thermodynamics in AdS Spacetime
In Fig. 1, upper left panel, we plot the metric function for Q=0.00,w=0.80Q=0.00,\ w=-0.80 and various values of the $a$ parameter. The metric tends to take smaller values as $a$ increases. In the upper right panel of the same figure, we plot the metric function for Q=0.00,a=0.80Q=0.00,\ a=0.80 and various values of the $w$ parameter. There is a different effect of $w$ on the small $(r\leq 1.0)$ and the large black holes. In the panel down left, we plot the metric function for Q=0.20,w=0.80Q=0.20,\ w=-0.80 and various values of the $a$ parameter. Finally, in the panel down right, we plot the metric function for Q=0.20,a=0.80Q=0.20,\ a=0.80 and various values of the $w$ parameter. As we can see, the existence of a non-zero charge $Q$ in the black hole tends to decrease the value of the metric function and to increase the $r_{+}$ value. The qualitative characteristics of the graphs are similar to the $Q=0.00$ cases.
@@
Figure 1: (a) Metric function $g\equiv f(r)$ versus $r$ for $Q=0.00,$ w=0.80w=-0.80 and $a=0.20,$ (blue curve, highest) $a=0.40$ (purple curve) $a=0.60$ (brown curve) and $a=0.80$ (cyan curve, lowest). (b) Metric function $g\equiv f(r)$ versus $r$ for $Q=0.00,$ $a=0.80$ and w=0.40,w=-0.40, (blue curve, highest on the right part) w=0.60w=-0.60 (purple curve) w=0.80w=-0.80 (brown curve) and w=1.00w=-1.00 (cyan curve, lowest on the right part). (c) Metric function $g\equiv f(r)$ versus $r$ for $Q=0.20,$ w=0.80w=-0.80 and $a=0.20,$ (blue curve, highest) $a=0.40$ (purple curve) $a=0.60$ (brown curve) and $a=0.80$ (cyan curve,lowest). (d) Metric function $g\equiv f(r)$ versus $r$ for $Q=0.20,$ $a=0.80$ and w=0.40,w=-0.40, (blue curve, highest on the right part) w=0.60w=-0.60 (purple curve) w=0.80w=-0.80 (brown curve) and w=1.00w=-1.00 (cyan curve, lowest on the right part).
@@
In Fig. 2, upper left panel, we plot the temperature versus $r_{h}$ for Q=0.00,w=0.80Q=0.00,\ w=-0.80 and various values of the $a$ parameter. The temperature decreases for increasing $a.$ Moreover the temperature is not a monotonic function and the small and large black holes respond differently to the quintessence. In the upper right panel of the same figure, we plot the temperature versus $r_{h}$ for Q=0.00,a=0.80Q=0.00,\ a=0.80 and various values of the $w$ parameter. As $|w|$ increases, the temperature is smaller for large $r_{h},$ while for modest $r_{h}$ it is larger. In the panel down left, we plot the temperature versus $r_{h}$ for Q=0.20,w=0.80Q=0.20,\ w=-0.80 and various values of the $a$ parameter. Finally, in the panel down right, we plot the temperature versus $r_{+}$ for Q=0.20,a=0.80Q=0.20,\ a=0.80 and various values of the $w$ parameter. For $Q=0.00,$ at some horizon size, the temperature has a minimum. The temperature of the black hole becomes smaller as $|w|$ increases from $\frac{1}{3}$ to $1$ for large enough black holes (big $r_{+}$), while it becomes larger for small $r_{+}.$ Thus large black holes radiate less for large $|w|,$ while small black holes radiate more. Thus large black holes seem to be longer-lived for large $|w|.$ For $Q>0$ the new feature is that the temperature presents a local minimum for small values of $r_{h},$ while it had a maximum for $Q=0.$
@@
Figure 2: (a) Temperature versus $r_{h}\equiv r_{+}$ for $Q=0.00,$ w=0.80w=-0.80 and $a=0.20,$ (blue curve, highest) $a=0.40$ (purple curve) $a=0.60$ (brown curve) and $a=0.80$ (cyan curve, lowest). (b) Temperature versus $r_{h}\equiv r_{+}$ for $Q=0.00,$ $a=0.80$ and w=0.40,w=-0.40, (blue curve, highest on the right part) w=0.60w=-0.60 (purple curve) w=0.80w=-0.80 (brown curve) and w=1.00w=-1.00 (cyan curve, lowest on the right part). (c) Temperature versus $r_{h}\equiv r_{+}$ for $Q=0.20,$ w=0.80w=-0.80 and $a=0.20,$ (blue curve, highest) $a=0.40$ (purple curve) $a=0.60$ (brown curve) and $a=0.80$ (cyan curve, lowest). (d) Temperature versus $r_{h}\equiv r_{+}$ for $Q=0.20,$ $a=0.80$ and w=0.40,w=-0.40, (blue curve, highest on the right part) w=0.60w=-0.60 (purple curve) w=0.80w=-0.80 (brown curve) and w=1.00w=-1.00 (cyan curve, lowest on the right part).
@@ IV Regge-Wheeler potentials
We concentrate specifically on the potential for scalar perturbations. The potential $V_{0}$ for zero spin (scalar) perturbations reads
@@
We start in Fig. 3, left panel, with the Regge-Wheeler potentials, when $a=0,$ so that quintessence plays no role at all; we set $M=0.05$ and let $Q$ take the values 0.00, 0.04, 0.070.00,\ 0.04,\ 0.07 and $0.10.$ It is obvious that, as $Q$ increases, the left barrier of the potential becomes smaller, and at $Q=0.10$ it disappears completely. In this case it is obvious that the potential cannot support bound states: any scalar excitation with $l=0$ will end into the black hole horizon. This characteristic is independent from the role of quintessence. We continue in the right panel with the case a=0.8,w=0.80a=0.8,\ w=-0.80 and $M=0.05$ and the same values for the charge $Q$ as before. The picture is qualitatively the same, but the presence of the quintessence tends to reduce the size of the potentials.
@@
Figure 3: Left panel: Potentials for a=0,l=0,M=0.05a=0,\ l=0,\ M=0.05 and four values of the charge: $Q=0.00$ (pink curve, highest), $Q=0.04$ (blue curve), $Q=0.07$ (khaki curve) and $Q=0.10$ (green curve, lowest). Right panel: Potentials for a=0.8,w=0.80,M=0.05a=0.8,\ w=-0.80,\ M=0.05 and $Q=0.00$ (pink curve, highest), $Q=0.04$ (blue curve), $Q=0.07$ (khaki curve) and $Q=0.10$ (green curve, lowest).
@@
We go ahead in Fig. 4 with the potentials at four different values of $w,$ namely 0.40,0.60,0.80,1.00,-0.40,\ -0.60,\ -0.80,\ -1.00, while l=0,M=0.05,l=0,\ M=0.05, for $Q=0.00$ (left panel) and $Q=0.04$ (right panel). We observe that, for both values of the charge, the potential for w=0.40w=-0.40 does not support bound states, but a left barrier gradually develops, as $|w|$ increases and bound states become possible.
@@
Figure 4: Left panel: Potentials for a=0.8,l=0,M=0.05,Q=0.00,a=0.8,\ l=0,\ M=0.05,\ Q=0.00, and four values for the parameter $w:$ w=0.40w=-0.40 (blue curve, no left barrier), w=0.60w=-0.60 (pink curve), w=0.80w=-0.80 (khaki curve) and w=1.00w=-1.00 (green curve, highest left barrier). Right panel: Same as in the left panel, but $Q=0.04.$
Fig. 5, left panel, depicts potentials for M=0.05,Q=0.00,w=0.60M=0.05,\ Q=0.00,\ w=-0.60 and l=0, 1, 2.l=0,\ 1,\ 2.. It clearly shows that the presence of angular momentum builds gradually a barrier on the left, so that more bound states may be formed. On the other hand, in the right panel one may inspect the potentials for M=0.05,Q=0.00,l=0,w=0.60M=0.05,\ Q=0.00,\ l=0,\ w=-0.60 and a=0.2, 0.4, 0.6, 0.8.a=0.2,\ 0.4,\ 0.6,\ 0.8. The potential decreases for increasing $a.$
@@
Figure 5: Left panel: Potentials for a=0.8,M=0.05,Q=0.00,w=0.60a=0.8,\ M=0.05,\ Q=0.00,\ w=-0.60 and three values for the angular momentum: $l=0$ (blue curve, lowest), $l=1$ (pink curve), and $l=2$ (green curve, highest). Right panel: Potentials for l=0,M=0.05,Q=0.00,w=0.60l=0,\ M=0.05,\ Q=0.00,\ w=-0.60 and four values for the parameter $a:$ $a=0.2$ (blue curve, highest), $a=0.4$ (pink curve), $a=0.6$ (khaki curve), and $a=0.8$ (green curve, lowest).
V Calculation of Bound States